OpenStudy (anonymous):

find the distance of a boat from a lighthouse if the lighthouse is 120 meters tall, and the angle of depression is 16degrees. sketch the problem situation. label the sketch with the given information. setup your right triangle trig function. box your solution. dont forget units

OpenStudy (jdoe0001):

 recall your SOH CAH TOA

OpenStudy (anonymous):

\[\tan(\theta)?\]

OpenStudy (jdoe0001):

yeap \(\bf tan(\theta)=\cfrac{opposite}{adjacent}\implies adjacent\cdot tan(\theta)=opposite \\ \quad \\ tan(16^o)=\cfrac{opposite}{120}\implies 120\cdot tan(16^o)=opposite\)